OpenStudy (anonymous):

help me.... √3:(-3125xy^(3))

OpenStudy (anonymous):

Can you put this in latex? All I see is a square root, then a 3, then a frowny face, then some random string of numbers/variables :D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

That way, I can help you :D

OpenStudy (anonymous):

\[\sqrt[3]{-3125xy ^{3}}\]

OpenStudy (anonymous):

COOL BEANS \[\sqrt[3]{-3125xy^3} = \sqrt[3]{-3125}*\sqrt[3]{x}*\sqrt[3]{y^3}\]\[5y \sqrt[3]{25x}\]

OpenStudy (anonymous):

Oops, put a negative sign at the front of that :D

OpenStudy (anonymous):

thats the answer that i got i just want to make sure i am doing this right :)

OpenStudy (anonymous):

Good work :D

OpenStudy (anonymous):

\[7/\sqrt[3]{49}\] what do you think about this?? I have an answer, but i want to see if you and i come up with the same

OpenStudy (anonymous):

if you don't mind can i ask you like other ten questions or so????